OpenStudy (anonymous):

Write the first derivative of the given function. f(x) = x^2[ln(x)]^18 I got: f'(x) = 2x(lnx)^18 + 18x(lnx)^17 Now part 2 asks, For which x-values does f(x) have horizontal tangents? I'm not getting it right so I need your help :) please and thank you!

OpenStudy (anonymous):

I think your first derivative is wrong. Try to do the product rule again. Remember that you will have to do the chain rule for (lnx)^18

OpenStudy (anonymous):

Oh god, you got the problem that I struggled for 1 week to solve... (they had us get the f13) anyways. x^2[ln(x)]^18 a = x^2 b = [ln(x)]^18 a' = 2x b' = [ln(x)]^17*(18/x) f' = 2x([ln(x)]^18) + (x^2)([ln(x)]^17*(18/x)) horizontal = 0 = f' 0=2x([ln(x)]^18) + (x^2)([ln(x)]^17*(18/x)) x = 1

OpenStudy (anonymous):

it's not 1. what does your derivative looks different from mine?

OpenStudy (anonymous):

ah wait, entered it wrong... 2 x log^17(x) (9+log(x)) x1 = 1 x2 = 1/e^9

OpenStudy (anonymous):

i realize that to find horizontal tangent you put f'(x) = 0. but it's so complicated. do you find the limit as it approaches 0. is there a method?

OpenStudy (cwrw238):

@kimmy the second part of your answer is incorrect its 18(lnx)^17 * 1/x * x^2 = 18x (lnx)^17 - excuse me - its correct!!

OpenStudy (cwrw238):

for horizontal tangents f'(x) = 0

OpenStudy (anonymous):

dreadslicer got the answer right? but how??

OpenStudy (anonymous):

Once you have the correct derivative, you need to set it equal to zero. Every solution will be a horizontal tangent.

OpenStudy (anonymous):

but using the derivative i provided (which is correct), how could you find a solution with ln? i only got x=1

OpenStudy (anonymous):

x can't be 0 because ln has to be greater than 0. that's why i'm not sure how he got 1/e^9